# CuriGS: Curriculum-Guided Gaussian Splatting for Sparse View Synthesis

### v1

@@ 1 Introduction
In recent years, 3D reconstruction has advanced rapidly, becoming a key enabling technology in fields such as virtual reality, digital twins, cultural heritage preservation, and intelligent manufacturing [yu2024sensors, tao2022digital, skublewska20223d]. Among various approaches [schonberger2016structure, wang2018pixel2mesh, qi2017pointnet], Neural Radiance Fields (NeRF) [mildenhall2021nerf] and its variants [barron2021mip, somraj2023vip, chen2021mvsnerf, pumarola2021d] have achieved impressive results in high-fidelity reconstruction and photorealistic view synthesis using implicit volumetric representations, albeit with slow optimization and heavy computational costs. Recently, 3D Gaussian Splatting (3DGS) [kerbl20233d] has emerged as a more efficient and stable explicit representation, enabling real-time rendering by modeling scenes as point-based Gaussians with covariance, color, and opacity attributes [qian20243dgs, liu2024citygaussian, wang2024view, mallick2024taming, zhang2025wheat3dgs, szymanowicz2024splatter]. Beyond representation efficiency, sparse view synthesis [long2022sparseneus, han2024binocular, cai2024structure, xu2024grm, charatan2024pixelsplat] has become a key research focus, aiming to reconstruct accurate geometry and realistic appearance from only a few input images. This setting is both practical and highly challenging, as limited viewpoints lead to supervision scarcity and severe overfitting.
In response to these challenges, recent studies have explored various strategies for enhancing 3D reconstruction under sparse-input conditions. Within the NeRF family, SparseNeRF [wang2023sparsenerf] introduces weak supervision by distilling depth-ordering priors, thereby constraining volumetric geometry and alleviating degradation under extremely limited viewpoints. FreeNeRF [yang2023freenerf] approaches the problem from a frequency and occlusion regularization perspective, imposing simple yet effective constraints on NeRF’s frequency spectrum and near-camera density distribution, which improves stability and generalization for few-shot rendering without external supervision.
In the context of 3DGS, FSGS [zhu2024fsgs] adapts 3DGS to sparse view synthesis through an efficient Gaussian expansion strategy, incorporating pretrained monocular depth estimation as a geometric prior to guide reconstruction. Similarly, DNGaussian [li2024dngaussian] exploits depth information by employing a hard and soft depth‑regularization scheme alongside a global‑local depth‑normalization strategy to reshape Gaussian primitives for more accurate geometry. Moreover, LoopSparseGS [bao2025loopsparsegs] proposes a loop-based framework that progressively densifies Gaussian initialization, aligns both scale-consistent and inconsistent depths for reliable geometry, and adopts sparse-aware sampling to mitigate oversized ellipsoid artifacts.
@@
Motivated by the above observations, we propose CuriGS, a curriculum-guided 3DGS framework designed for sparse view 3D reconstruction. The key idea of CuriGS is the introduction of student views—pseudo-views generated around real cameras (teacher) with controllable perturbation magnitudes. During training, a curriculum schedule progressively unlocks student groups with larger perturbations, allowing the model to gradually adapt from locally consistent to more diverse viewpoints. At each iteration, a subset of student views is randomly sampled from the currently active group and optimized through depth-correlation and co-regularization constraints, which enforce geometric consistency while mitigating reconstruction errors. Meanwhile, CuriGS maintains the best-performing student for each teacher and perturbation group, evaluated via a multi-signal metric combining structural similarity (SSIM), perceptual similarity (LPIPS), and a no-reference image-quality score [agnolucci2024qualityaware]. Students that exceed a predefined quality threshold are periodically promoted into the training set, effectively augmenting sparse supervision with reliable pseudo-views. This design effectively enriches the supervision signal under sparse-view conditions, improving both geometric fidelity and rendering realism while enhancing the model’s generalization to unseen viewpoints. The main contributions of this work are as follows:
@@ 2.1 NeRF-based Sparse View Synthesis
While NeRF achieves photorealistic results under dense multi-view supervision, its optimization becomes severely ill-posed in sparse-view settings, often leading to geometry collapse, texture ambiguity, and overfitting to the observed views. To alleviate these issues, several methods [jain2021putting, somraj2023vip, deng2022depth, niemeyer2022regnerf, somraj2024simple, wang2023sparsenerf, yang2023freenerf] have been proposed to adapt NeRF-style representations for few-shot reconstruction.
SparseNeRF [wang2023sparsenerf] introduces weak geometric supervision by distilling depth-ordering priors from pretrained monocular depth estimators or coarse sensor depth. By injecting such constraints into the volumetric optimization, SparseNeRF reduces geometric ambiguity and mitigates depth collapse when only a handful of views are available. However, its performance heavily depends on the quality and domain alignment of the external depth priors, which may mislead optimization when the monocular estimator is biased or out-of-domain. FreeNeRF [yang2023freenerf] approaches the problem from a spectral and occlusion-regularization perspective, imposing frequency and near-camera density constraints on the learned radiance field to suppress unsupported high-frequency components and spurious near-field densities that commonly arise under sparse supervision. This family of regularizers is appealing for its simplicity and independence from explicit external priors, but it requires careful hyperparameter tuning and may attenuate genuine high-frequency surface details in richly textured scenes. PixelNeRF [yu2021pixelnerf] extends NeRF into a conditional formulation by leveraging learned multi-view feature fusion. Through large-scale multi-scene pretraining, it improves few-shot generalization and cross-scene adaptability, yet still struggles with limited fidelity and unstable geometry when applied to novel domains that diverge from the training distribution. Despite these advances, sparse-view NeRFs are fundamentally constrained by their implicit volumetric representation. Consequently, achieving stable geometry and consistent appearance reconstruction under extremely sparse inputs remains a significant challenge, motivating the development of alternative representations and training paradigms that can inherently cope with data scarcity.
@@
Compared with NeRF, 3DGS offers superior runtime efficiency, interpretable geometric primitives, and direct manipulability, making it particularly appealing for interactive visualization and efficient optimization. Motivated by these strengths, several studies [li2024dngaussian, zhu2024fsgs, zhang2024cor, zheng2025nexusgs, park2025dropgaussian, paliwal2024coherentgs, liu2026sv, zhang2026usgs] have explored its potential for sparse view synthesis tasks with most achieving notably better performance than NeRF-based methods.
CoR-GS [zhang2024cor] introduces an ensemble-based strategy in which multiple 3DGS instances are trained jointly, using mutual inconsistencies in both point-level geometry and rendered appearance to identify and suppress unreliable splats through co-pruning. This co-regularization effectively reduces reconstruction artifacts without relying on explicit ground-truth geometry. However, the method increases training complexity and computational overhead, and its effectiveness diminishes in highly symmetric or extremely sparse scenes where ensemble disagreement becomes ambiguous. NexusGS [zheng2025nexusgs] introduces an epipolar-depth–guided framework that integrates geometric priors to guide Gaussian densification, pruning, and blending. However, it relies on additional optical flow information and performs poorly when training views have low overlap, which limits its applicability. DropGaussian [park2025dropgaussian] proposes a lightweight structural regularization strategy by randomly deactivating subsets of Gaussians during training. This dropout-like mechanism redistributes gradients toward under-optimized splats, mitigating overfitting and improving generalization to unseen views. Unlike methods that rely on external priors or ensemble models, DropGaussian is simple and computationally efficient, though its stochastic nature may lead to unstable convergence or suboptimal results in extremely sparse or unstructured scenes.
Beyond the representative works discussed above, other 3DGS-based sparse-view methods [tang2025spars3r, xiong2025sparsegs, wan2025s2gaussian, bao2025loopsparsegs] have made progress by incorporating semantic guidance, Gaussian super-resolution, or confidence-weighted sparse enhancements to stabilize optimization. Nevertheless, these approaches fundamentally build upon stronger regularization or auxiliary supervision and do not address the key challenge of sparse view reconstruction, namely the intrinsic scarcity of training viewpoints, thus remaining methodologically constrained. In contrast, we directly tackle the fundamental limitation of sparse training supervision by expanding the available viewpoints through a curriculum-guided pseudo-view learning strategy.
@@ 3.3 Student View Evaluation and Promotion
At each training iteration, for each teacher view $\mathcal{C}_{t}$, one student view $\mathcal{C}_{s}^{\sigma_{\text{active}},j}\in\mathcal{P}(\sigma_{\text{active}})$ is randomly sampled from the pool of student views corresponding to the currently active perturbation level $\sigma_{\text{active}}$. The corresponding rendered image is then compared against the teacher’s reference image through a composite multi-signal metric. This metric is designed to provide an objective evaluation of student views in the absence of ground-truth references, jointly accounting for structural similarity (SSIM) to assess spatial fidelity, perceptual distance (LPIPS) to measure feature-level realism, and a no-reference image-quality measure [agnolucci2024qualityaware] to capture global perceptual quality. For each $(\mathcal{C}_{t},\sigma_{i})$ pair, we maintain the best-performing student, defined as the candidate with the lowest evaluation loss up to the current iteration.
@@
3.4 Losses and Training Objective
Our overall training objective consists of three complementary parts: (1) a photometric loss computed over all current training views, including the promoted student views; (2) a photometric loss on the original ground-truth views; and (3) the student-view regularization term.
Training and Ground-Truth Views.
While newly promoted student views effectively expand the sparse training distribution, their pseudo-supervision may inevitably introduce uncertainty. To prevent potential drift and maintain stable optimization, we additionally preserve explicit supervision on the original teacher views throughout training. This dual-level formulation allows the network to continuously refine its geometry and appearance while anchoring the learning process to reliable real observations. We employ a hybrid photometric loss that combines $\mathcal{L}_{\text{1}}$ and D-SSIM:
$\mathcal{L}_{\text{recon}}=(1-\lambda_{s})\mathcal{L}_{\text{1}}+\lambda_{s}\mathcal{L}_{\text{ssim}},$ (5)
Student Views.
Since student views are not real camera observations, their rendered outputs lack ground-truth images for direct supervision. Therefore, we introduce two pseudo-supervision schemes to optimize the learning of student views. The first scheme is based on depth-correlation. Specifically, we employ a pretrained monocular depth estimation model [ranftl2021vision] to predict a pseudo-depth map from the rendered image of a student view. This pseudo-depth is then compared with the depth map directly output by the 3DGS rasterization. To quantify their structural alignment, we adopt the Pearson correlation to formulate a relaxed relative loss. The formula:
$\mathcal{L}_{\text{depth}}^{s}=1-\frac{\mathrm{Cov}(D_{r},D_{p})}{\sigma(D_{r})\,\sigma(D_{p})},$ (6)
The second scheme relies on a dual-model consistency constraint [zhang2024cor]. During training, we maintain two 3DGS models with independent parameters, rendering the same student view in parallel. The photometric loss between their rendered results serves as a self-supervised constraint, enforcing cross-model visual consistency and reducing uncertainty introduced by view perturbations. We define the final student loss as:
$\mathcal{L}_{\text{stu}}=\lambda_{d}\,\mathcal{L}_{\text{depth}}^{s}+\lambda_{p}\,\mathcal{L}_{\text{photo}}^{s}.$ (7)
Overall Objective.
The overall optimization objective is formulated as:
$\mathcal{L}=\lambda_{\text{1}}\,\mathcal{L}_{\text{train}}+\lambda_{\text{2}}\,\mathcal{L}_{\text{gt}}+\lambda_{\text{3}}\,\mathcal{L}_{\text{stu}},$ (8)
@@ 4 Experiments
RegNeRF[niemeyer2022regnerf] 19.08 0.587 0.336 22.19 0.643 0.335 18.89 0.745 0.190
SparseNeRF[wang2023sparsenerf] 19.86 0.624 0.328 22.85 0.693 0.315 19.55 0.769 0.201
FreeNeRF[yang2023freenerf] 19.63 0.612 0.308 22.78 0.689 0.323 19.92 0.787 0.182
3DGS[kerbl20233d] 18.54 0.588 0.272 21.71 0.672 0.248 17.65 0.816 0.146
DNGaussian[li2024dngaussian] 19.12 0.591 0.294 18.06 0.423 0.584 18.91 0.790 0.176
FSGS[zhu2024fsgs] 20.43 0.682 0.248 23.40 0.733 0.238 17.14 0.818 0.162
CoR-GS[zhang2024cor] 20.45 0.712 0.196 23.55 0.727 0.226 19.21 0.853 0.119
DropGaussian[park2025dropgaussian] 20.76 0.713 0.200 23.66 0.747 0.233 19.31 0.857 0.142
NexusGS[zheng2025nexusgs] 21.00 0.730 0.179 23.86 0.753 0.206 20.21 0.869 0.102
LoopSparseGS[bao2025loopsparsegs] 20.85 0.717 0.205 24.09 0.755 0.226 20.68 0.856 0.125
Ours 21.10 0.732 0.193 24.10 0.759 0.205 20.35 0.870 0.131
@@
4.1 Datasets & Baselines
Datasets.
We evaluated our method on three widely adopted datasets: LLFF [mildenhall2019local], DTU [jensen2014large], and Mip-NeRF360 [barron2022mip]. Across all datasets, we assume that the camera parameters are known. All experimental settings, including data splits and image resolutions, are kept consistent with previous studies [zhu2024fsgs, zheng2025nexusgs, park2025dropgaussian] to ensure fair comparison.
Baselines.
We compared our approach with a representative set of state-of-the-art methods from two categories, including NeRF-based approaches [niemeyer2022regnerf, wang2023sparsenerf, yang2023freenerf] and recent 3DGS-based methods [li2024dngaussian, zhu2024fsgs, zhang2024cor, zheng2025nexusgs, park2025dropgaussian]. For quantitative evaluation, we reported the standard image space metrics PSNR, SSIM, and LPIPS computed on held-out test views.

### v2

@@ 1 Introduction
In recent years, 3D reconstruction has advanced rapidly, becoming a key enabling technology in fields such as virtual reality, digital twins, cultural heritage preservation, and intelligent manufacturing [45, 34, 29, 8]. Among various approaches [28, 38, 25], Neural Radiance Fields (NeRF) [20] and its variants [3, 31, 7, 24] have achieved impressive results in high-fidelity reconstruction and photorealistic view synthesis using implicit volumetric representations, albeit with slow optimization and heavy computational costs. Recently, 3D Gaussian Splatting (3DGS) [13] has emerged as a more efficient and stable explicit representation, enabling real-time rendering by modeling scenes as point-based Gaussians with covariance, color, and opacity attributes [26, 16, 40, 18, 46, 32, 37, 39]. Beyond representation efficiency, sparse view synthesis [17, 10, 5, 42, 6] has become a key research focus, aiming to reconstruct accurate geometry and realistic appearance from only a few input images. This setting is both practical and highly challenging, as limited viewpoints lead to supervision scarcity and severe overfitting.
In response to these challenges, recent studies have explored various strategies for enhancing 3D reconstruction under sparse-input conditions. Within the NeRF family, SparseNeRF [36] introduces weak supervision by distilling depth-ordering priors, thereby constraining volumetric geometry and alleviating degradation under extremely limited viewpoints. FreeNeRF [43] approaches the problem from a frequency and occlusion regularization perspective, imposing simple yet effective constraints on NeRF’s frequency spectrum and near-camera density distribution, which improves stability and generalization for few-shot rendering without external supervision.
In the context of 3DGS, FSGS [50] adapts 3DGS to sparse view synthesis through an efficient Gaussian expansion strategy, incorporating pretrained monocular depth estimation as a geometric prior to guide reconstruction. Similarly, DNGaussian [14] exploits depth information by employing a hard and soft depth‑regularization scheme alongside a global‑local depth‑normalization strategy to reshape Gaussian primitives for more accurate geometry. Moreover, LoopSparseGS [2] proposes a loop-based framework that progressively densifies Gaussian initialization, aligns both scale-consistent and inconsistent depths for reliable geometry, and adopts sparse-aware sampling to mitigate oversized ellipsoid artifacts.
@@
Motivated by the above observations, we propose CuriGS, a curriculum-guided 3DGS framework designed for sparse view 3D reconstruction. The key idea of CuriGS is the introduction of student views—pseudo-views generated around real cameras (teacher) with controllable perturbation magnitudes. During training, a curriculum schedule progressively unlocks student groups with larger perturbations, allowing the model to gradually adapt from locally consistent to more diverse viewpoints. At each iteration, a subset of student views is randomly sampled from the currently active group and optimized through depth-correlation and co-regularization constraints, which enforce geometric consistency while mitigating reconstruction errors. Meanwhile, CuriGS maintains the best-performing student for each teacher and perturbation group, evaluated via a multi-signal metric combining structural similarity (SSIM), perceptual similarity (LPIPS), and a no-reference image-quality score [1]. Students that exceed a predefined quality threshold are periodically promoted into the training set, effectively augmenting sparse supervision with reliable pseudo-views. As illustrated in Fig. 1, this design effectively enriches the supervision signal under sparse-view conditions, improving both geometric fidelity and rendering realism while enhancing the model’s generalization to unseen viewpoints. The main contributions of this work are as follows:
@@ 2.1 NeRF-based Sparse View Synthesis
While NeRF achieves photorealistic results under dense multi-view supervision, its optimization becomes severely ill-posed in sparse-view settings, often leading to geometry collapse, texture ambiguity, and overfitting to the observed views. To alleviate these issues, several methods [11, 31, 9, 21, 30, 36, 43] have been proposed to adapt NeRF-style representations for few-shot reconstruction.
SparseNeRF [36] introduces weak geometric supervision by distilling depth-ordering priors from pretrained monocular depth estimators or coarse sensor depth. By injecting such constraints into the volumetric optimization, SparseNeRF reduces geometric ambiguity and mitigates depth collapse when only a handful of views are available. However, its performance heavily depends on the quality and domain alignment of the external depth priors, which may mislead optimization when the monocular estimator is biased or out-of-domain. FreeNeRF [43] approaches the problem from a spectral and occlusion-regularization perspective, imposing frequency and near-camera density constraints on the learned radiance field to suppress unsupported high-frequency components and spurious near-field densities that commonly arise under sparse supervision. This family of regularizers is appealing for its simplicity and independence from explicit external priors, but it requires careful hyperparameter tuning and may attenuate genuine high-frequency surface details in richly textured scenes. PixelNeRF [44] extends NeRF into a conditional formulation by leveraging learned multi-view feature fusion. Through large-scale multi-scene pretraining, it improves few-shot generalization and cross-scene adaptability, yet still struggles with limited fidelity and unstable geometry when applied to novel domains that diverge from the training distribution. Despite these advances, sparse-view NeRFs are fundamentally constrained by their implicit volumetric representation. Consequently, achieving stable geometry and consistent appearance reconstruction under extremely sparse inputs remains a significant challenge, motivating the development of alternative representations and training paradigms that can inherently cope with data scarcity.
@@
Compared with NeRF, 3DGS offers superior runtime efficiency, interpretable geometric primitives, and direct manipulability, making it particularly appealing for interactive visualization and efficient optimization. Motivated by these strengths, several studies [14, 50, 47, 49, 23, 22, 15, 48] have explored its potential for sparse view synthesis tasks with most achieving notably better performance than NeRF-based methods.
CoR-GS [47] introduces an ensemble-based strategy in which multiple 3DGS instances are trained jointly, using mutual inconsistencies in both point-level geometry and rendered appearance to identify and suppress unreliable splats through co-pruning. This co-regularization effectively reduces reconstruction artifacts without relying on explicit ground-truth geometry. However, the method increases training complexity and computational overhead, and its effectiveness diminishes in highly symmetric or extremely sparse scenes where ensemble disagreement becomes ambiguous. NexusGS [49] introduces an epipolar-depth–guided framework that integrates geometric priors to guide Gaussian densification, pruning, and blending. However, it relies on additional optical flow information and performs poorly when training views have low overlap, which limits its applicability. DropGaussian [23] proposes a lightweight structural regularization strategy by randomly deactivating subsets of Gaussians during training. This dropout-like mechanism redistributes gradients toward under-optimized splats, mitigating overfitting and improving generalization to unseen views. Unlike methods that rely on external priors or ensemble models, DropGaussian is simple and computationally efficient, though its stochastic nature may lead to unstable convergence or suboptimal results in extremely sparse or unstructured scenes.
Beyond the representative works discussed above, other 3DGS-based sparse-view methods [33, 41, 35, 2] have made progress by incorporating semantic guidance, Gaussian super-resolution, or confidence-weighted sparse enhancements to stabilize optimization. Nevertheless, these approaches fundamentally build upon stronger regularization or auxiliary supervision and do not address the key challenge of sparse view reconstruction, namely the intrinsic scarcity of training viewpoints, thus remaining methodologically constrained. In contrast, we directly tackle the fundamental limitation of sparse training supervision by expanding the available viewpoints through a curriculum-guided pseudo-view learning strategy.
@@ 3.3 Student View Evaluation and Promotion
At each training iteration, for each teacher view $\mathcal{C}_{t}$, one student view $\mathcal{C}_{s}^{\sigma_{\text{active}},j}\in\mathcal{P}(\sigma_{\text{active}})$ is randomly sampled from the pool of student views corresponding to the currently active perturbation level $\sigma_{\text{active}}$. The corresponding rendered image is then compared against the teacher’s reference image through a composite multi-signal metric. This metric is designed to provide an objective evaluation of student views in the absence of ground-truth references, jointly accounting for structural similarity (SSIM) to assess spatial fidelity, perceptual distance (LPIPS) to measure feature-level realism, and a no-reference image-quality measure [1] to capture global perceptual quality. For each $(\mathcal{C}_{t},\sigma_{i})$ pair, we maintain the best-performing student, defined as the candidate with the lowest evaluation loss up to the current iteration.
@@
3.4 Optimization Objective
To ensure robust reconstruction that benefits from dataset expansion while maintaining fidelity to the ground truth, we formulate our objective function with three distinct components: (1) a dynamic reconstruction loss computed over the current training set (comprising both original and promoted student views) to drive the primary optimization; (2) an anchor loss specifically enforced on the fixed original ground-truth views to prevent semantic drift; and (3) student-view geometric regularizers including depth-correlation and dual-model consistency to constrain the structure. The total loss $\mathcal{L}_{\text{total}}$ is defined as:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{train}}+\lambda_{\text{drift}}\,\mathcal{L}_{\text{anchor}}+\lambda_{\text{reg}}\,\mathcal{L}_{\text{reg}}$ (5)
Dynamic Training Loss.
This term drives the primary optimization of the 3D Gaussian field. It is computed over the current training set $\mathcal{V}_{\text{train}}$, which is dynamically updated throughout the curriculum. Initially, $\mathcal{V}_{\text{train}}$ contains only the sparse input views. As training progresses, high-confidence student views (pseudo-views that pass our quality curriculum) are added to $\mathcal{V}_{\text{train}}$.
[…]
where $\mathcal{B}_{\text{train}}$ is a batch sampled from $\mathcal{V}_{\text{train}}$, and $\mathcal{L}_{\text{photo}}$ combines L1 and D-SSIM loss. This allows the model to continuously learn from the densified view coverage.
Anchor Loss.
As the number of synthetic views in $\mathcal{V}_{\text{train}}$ grows, there is a risk that the optimization may drift or overfit to artifacts in the generated data. To prevent this, we enforce a strict consistency constraint on the fixed set of original input views $\mathcal{V}_{\text{t}}$ (the teacher views). This set remains static throughout training and acts as a geometric anchor. To maintain efficiency, we randomly sample one anchor view $I_{\text{anc}}\sim\mathcal{V}_{\text{t}}$ at each iteration:
$\mathcal{L}_{\text{anchor}}=\mathcal{L}_{\text{photo}}(I_{\text{anc}},\hat{I}_{\text{anc}})$ (7)
By explicitly re-visiting the ground truth at every step, we ensure that the geometric optimization remains anchored to the reliable observations, preventing semantic drift without incurring the computational cost of rendering all original views.
Student Regularization Loss.
Since student views are synthetic observations sampled from the curriculum, they lack ground-truth pixel data for direct supervision. Optimizing these views without constraints can lead to hallucinated geometry or floaters. Therefore, we introduce two pseudo-supervision schemes to regularize the structure of student views. The first scheme is based on depth-correlation. Specifically, we leverage a pretrained monocular depth estimation model [27] to extract a proxy depth map $D_{\text{proxy}}$ from the rendered student image. This proxy depth is then compared with the metric depth $D_{\text{render}}$ directly output by the 3DGS rasterization via the Pearson correlation coefficient, ensuring that the 3DGS geometry structurally aligns with the visual cues without enforcing an erroneous absolute scale:
$\mathcal{L}_{\text{depth}}=1-\frac{\text{Cov}(D_{\text{render}},D_{\text{proxy}})}{\sigma_{\text{render}}\sigma_{\text{proxy}}}$ (8)
where Cov denotes covariance and $\sigma$ denotes standard deviation. The second scheme employs a dual-model consistency constraint. Inspired by recent co-regularization strategies [47], we leverage the insight that valid geometry tends to be consistent across independently trained models, whereas artifacts are typically stochastic and variable. Consequently, we maintain two independently initialized models, $\mathcal{M}_{A}$ and $\mathcal{M}_{B}$, and enforce photometric consistency between their renderings of the same student view, encouraging the models to reach a consensus on the underlying scene structure:
$\mathcal{L}_{\text{co}}=\|I_{\text{render}}^{A}-I_{\text{render}}^{B}\|_{2}^{2}$ (9)
The final regularization term is the weighted sum:
$\mathcal{L}_{\text{reg}}=\lambda_{\text{d}}\mathcal{L}_{\text{depth}}+\lambda_{\text{c}}\mathcal{L}_{\text{co}}$ (10)
@@ 4 Experiments
RegNeRF[21] 19.08 0.587 0.336 22.19 0.643 0.335 18.89 0.745 0.190
SparseNeRF[36] 19.86 0.624 0.328 22.85 0.693 0.315 19.55 0.769 0.201
FreeNeRF[43] 19.63 0.612 0.308 22.78 0.689 0.323 19.92 0.787 0.182
3DGS[13] 18.54 0.588 0.272 21.71 0.672 0.248 17.65 0.816 0.146
DNGaussian[14] 19.12 0.591 0.294 18.06 0.423 0.584 18.91 0.790 0.176
FSGS[50] 20.43 0.682 0.248 23.40 0.733 0.238 17.14 0.818 0.162
CoR-GS[47] 20.45 0.712 0.196 23.55 0.727 0.226 19.21 0.853 0.119
DropGaussian[23] 20.76 0.713 0.200 23.66 0.747 0.233 19.31 0.857 0.142
NexusGS[49] 21.00 0.730 0.179 23.86 0.753 0.206 20.21 0.869 0.102
LoopSparseGS[2] 20.85 0.717 0.205 24.09 0.755 0.226 20.68 0.856 0.125
Ours 21.10 0.732 0.193 24.21 0.761 0.202 20.45 0.873 0.129
@@
4.1 Datasets
To comprehensively evaluate the robustness and generalization ability of the proposed CuriGS framework, we conduct extensive experiments on three diverse and widely adopted benchmarks: LLFF [19], MipNeRF-360 [4], and DTU [12]. These datasets encompass a variety of challenging scenarios, ranging from controlled object-centric captures to complex unbounded real-world environments. To ensure a strictly fair comparison, all data splits, image resolutions, and preprocessing steps are kept consistent with prior sparse-view studies [50, 49, 23, 14, 47].
LLFF.
The LLFF dataset [19] consists of complex, forward-facing real-world scenes captured by handheld cameras. Following the standard evaluation protocol, we designate every 8th image as the test set and utilize the remainder as the training pool. To simulate extreme sparse-view conditions, we uniformly subsample a highly restricted number of views from the training pool to serve as our input. All images are downsampled by a factor of 8 to maintain computational efficiency while aligning with baseline configurations.
MipNeRF-360.
The MipNeRF-360 dataset [4] presents highly challenging unbounded indoor and outdoor scenes with intricate background details and 360-degree camera trajectories. This dataset thoroughly tests the model’s capacity to handle scale variations and unbounded backgrounds under sparse supervision. Similar to the LLFF protocol, we reserve every 8th image for testing and uniformly sample a sparse subset from the remaining images for training. The image resolution is also downsampled by a factor of 8.
DTU.
The DTU dataset [12] features object-centric scenes captured under strictly controlled laboratory conditions with precise camera poses. For the DTU dataset, we focus our evaluation on 15 representative scenes, specifically scans 8, 21, 30, 31, 34, 38, 40, 41, 45, 55, 63, 82, 103, 110, and 114. For the extreme sparse-view setting, we specifically utilize views 25, 22, and 28 to construct a 3-view training set, reserving the remaining 25 views for novel view evaluation. All images are downsampled by a factor of 4. Furthermore, to strictly focus the optimization and evaluation on the geometric and photometric fidelity of the target objects, binary masks are applied to isolate and remove the background regions.
4.2 Implementation Details
To handle the varying complexity and scene types across different datasets, we adopt dataset-specific training configurations as detailed below.
LLFF.
For the LLFF scenes, the model is trained for 30,000 iterations. The curriculum-based student sampling phase is active from iteration 3,000 to 24,000. In this phase, we generate pseudo-student views using perturbation levels $\sigma\in\{1,2,3,4,5,6,7,8,9,10\}$, with 5 students per level. The quality threshold for selecting reliable student views is set to 0.4.
MipNeRF-360.
In the MipNeRF-360 dataset, training generally runs for 30,000 iterations, with the curriculum activation scheduled between iterations 7,000 and 27,000. For more complex scenes such as bicycle and stump, we extend the training to 33,000 iterations and the curriculum phase from 10,000 to 30,000. Student views are generated using perturbation levels $\sigma\in\{2,4,6,8,10\}$, with 3 students per level and a quality threshold of 0.45.
DTU.
For the DTU dataset, the model is optimized for 13,000 iterations, with curriculum guidance applied between 2,000 and 12,000 iterations. We utilize perturbation levels $\sigma\in\{1,2,3,4,5\}$ and sample 10 students per level. The quality threshold is set between 0.45 and 0.55. To maintain object-centric consistency, we apply an adaptive masking strategy for the sampled student views. For a specific student view $\mathcal{C}_{s}$, its background mask $M_{s}$ is determined by the color statistics of the background regions $\mathcal{B}_{t}$ from its corresponding teacher view $\mathcal{C}_{t}$. Let $\boldsymbol{\mu}_{s}$ and $\boldsymbol{\sigma}_{s}$ be the mean and standard deviation of pixel colors in $\mathcal{C}_{s}$ at coordinates corresponding to $\mathcal{B}_{t}$. A pixel $p$ in the student view is classified as background if:
$M_{s}(p)=\begin{cases}1,&\text{if }\left\lVert\mathcal{C}_{s}(p)-\boldsymbol{\mu}_{s}\right\rVert<\tau\,\boldsymbol{\sigma}_{s},\\ 0,&\text{otherwise},\end{cases}$ (11)
where $\tau$ is a tolerance threshold. This mechanism ensures that the optimization remains focused on the target object by robustly suppressing background noise, even when geometric perturbations are introduced during the student camera sampling process.
4.3 Baselines
We compared our approach with a representative set of state-of-the-art methods from two categories, including NeRF-based approaches [21, 36, 43] and recent 3DGS-based methods [14, 50, 47, 49, 23]. For quantitative evaluation, we reported the standard image space metrics PSNR, SSIM, and LPIPS computed on held-out test views.